OpenStudy (anonymous):

If θ = 68° and the side opposite θ has length 114.18, how long is the hypotenuse?

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

\[\sin\theta=\frac{\text{opposite side}}{\text{hypotenuse}}\]

OpenStudy (unklerhaukus):

rearrange the definition for the sine of an angle, to solve for the hypotenuse , then just substitute in the values